# Strand, London WC2R 2LS, UK

### v1

@@ A.3 Variation of the Maxwell Action
In our conventions (9) the Maxwell action reads
$\mathcal{I}_{\text{M}}=\frac{1}{64\pi G_{N}}\int_{\mathcal{M}}d^{d+1}x\sqrt{g}F^{\mu\nu}F_{\mu\nu}\ .$ (147)
@@
$\mathcal{I}_{\text{M}}=\frac{1}{32\pi G_{N}}\int_{\mathcal{M}}F\wedge\star F\ .$ (148)
@@
$\delta\mathcal{I}_{\text{M}}=\frac{1}{16\pi G_{N}}\int_{\mathcal{M}}d\delta A\wedge\star F=\frac{1}{16\pi G_{N}}\left(\int_{\partial\mathcal{M}}\delta A\wedge\star F-\int_{\mathcal{M}}\delta A\wedge d\star F\right)\ ,$ (149)
@@
$\delta\mathcal{I}_{\text{M}}\hat{=}\frac{1}{16\pi G_{N}}\int_{\partial\mathcal{M}}\delta A\wedge\star F\ ,$ (151)
@@
$\delta\mathcal{I}_{\text{M}}\hat{=}\frac{1}{16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}A\wedge\star F\right)-\frac{1}{16\pi G_{N}}\int_{\partial\mathcal{M}}A\wedge\delta(\star F)\ ,$ (152)
@@
$\mathcal{I}_{\text{M}}^{(\partial\mathcal{A})}=-\frac{1}{16\pi G_{N}}\int_{\partial\mathcal{M}}A\wedge\star F\equiv-\frac{1}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}F^{\mu\nu}n_{\mu}A_{\nu}\ .$ (153)
then from the remaining term of eq. (152), the required boundary condition for a good variational problem is $\delta(\sqrt{h}F^{\mu\nu}n_{\mu})|_{\partial\mathcal{M}}=0$. This corresponds to fixing the electric charge, since
@@
From the above derivation it is also evident that boundary conditions on the metric and the gauge field can be imposed independently.
@@ Appendix B Local and Covariant Boundary Terms for CBC
The purpose of this appendix is to demonstrate that there does not exist a local and covariant boundary term which yields a well-defined variational principle for the bulk action (9) subject to the CBC (3).
To begin, note that after taking into account the Gauss-Codazzi equations, there exist nine independent local and covariant boundary terms one can write down, whose integrands take the form Bunch1981 (again curvature tensors with overhead tilde are evaluated for the boundary metric)
@@
where $\mathcal{L}_{n}$ is the Lie derivative with respect to the normal $n$ to the boundary $\partial\mathcal{M}$. As noted e.g., in ref Bunch1981 , the variations of $D^{i}D_{i}\mathcal{K},\,D_{i}D_{j}\mathcal{K}^{ij}$ yield third order derivatives of the metric variation, and there is no linear combination of the two or some contribution from the variations of any of the other of the seven terms of eq. (155) that can cancel them. Hence, these cannot be included in the boundary term. A similar argument shows that the variations of the last two terms in eq. (155) cannot be canceled either.
@@
We would now like to examine whether there exists a combination of the coefficients $c_{i}$, such that the variation of eq. (156) cancels the boundary pieces of the on-shell variation of the bulk E(M)GB action after imposing CBC (3). Since, for CBC, we fix the conformal class of the induced metric, it will be convenient to decompose the various variations in terms of the variation of the conformal class and the variation of the Weyl factor. To accomplish that, we recall that for two $d$-dimensional metrics that are related via a Weyl rescaling $\tilde{h}_{ij}=\Omega^{2}(y^{m})h_{ij}$, the extrinsic curvature, Ricci tensor and Ricci scalar are related by (see e.g., appendix D in Wald’s book Wald:1984rg )
$\begin{cases}&\tilde{\mathcal{K}}_{ij}=\Omega\big(\mathcal{K}_{ij}+h_{ij}n^{k}D_{k}\log{\Omega}\big)\ ,\\ &\begin{aligned} \tilde{R}_{ij}=R_{ij}&-(d-2)D_{i}D_{j}\log{\Omega}-h_{ij}h^{pq}D_{p}D_{q}\log{\Omega}\\ &+(d-2)D_{i}\log{\Omega}D_{j}\log{\Omega}-(d-2)h_{ij}h^{pq}D_{p}\log{\Omega}D_{q}\log{\Omega}\end{aligned}\ ,\\ &\tilde{R}=\Omega^{-2}\Big(R-2(d-1)h^{ij}D_{i}D_{j}\log{\Omega}-(d-1)(d-2)h^{ij}D_{i}\log{\Omega}D_{j}\log{\Omega}\Big)\ ,\end{cases}$ (157)
@@
$+h^{-\frac{1}{d}}\mathcal{K}D^{i}\log{\sqrt{h}}D_{i}\log{\sqrt{h}}\big)\Bigg)\delta\sqrt{h}\Bigg]$
@@ Appendix C On-Shell Actions for EMGB Solutions
In this appendix, we briefly describe how to compute the on-shell actions (9) for the solutions  (17). In computing the on-shell actions, the following curvature quantities for the general metric of the form (15) will be useful Deser:2005pc ; Haroon:2020vpr :
@@
where we defined $\lambda=d-1,\ a(r)=f^{\prime}(r)/2\sqrt{f(r)}\ \text{and}\ b(r)=\sqrt{f(r)}/r$. Moreover, the trace of the EGB field equations is (recall the definition of $\mathcal{E}_{\mu\nu}$ from eq. (117) and also that $\mathcal{L}_{\text{GB}}=R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}-4R^{\mu\nu}R_{\mu\nu}+R^{2}$):
@@
which has to be zero on-shell, and thus we learn:
@@
Using this, we can re-express the bulk term of the action (9) as:
@@
since the $r$ integral reduces to a boundary term by successive integrations by parts:
@@
A useful identity is the following (here $q=0$):
@@
Using eq. (165), the bulk on-shell action simplifies to:
@@
Therefore, the GB boundary term (119) can be collectively expressed as:
@@
while the corresponding GR boundary term (96) reads:
@@
Putting eqs. (166), (168), (169) together we have the following full on-shell action (here we specialize to $k=a_{d}=b_{d}=1$ and $q=0$):
@@ Appendix D Flat EMGB black hole
We consider first the case of an asymptotically flat BD black hole with spherical horizon ($q=0,\Lambda=0,k=1$ in eq. (17)), with line element:
@@
The boundary is at $r=e^{\boldsymbol{\omega}}\mathfrak{r}$ and we consider thermodynamics for the region between the black hole horizon at $r=r_{h}$ and the boundary $\partial\mathcal{M}$. As in (A)dS, inverting the boundary data eqs. (12) for the metric (171) to obtain $r_{h},e^{\boldsymbol{\omega}}\mathfrak{r}$ as a function of $\{\tilde{\beta},\mathcal{P}\}$ is not possible analytically, even for the simplest case of $d=4$. Nonetheless, perturbatively in the GB coupling, we get (recall that $y=r_{h}/e^{\boldsymbol{\omega}}\mathfrak{r}$):
@@
Due to the absence of additional scales in flat space, the high-temperature limit corresponds to sending $y\rightarrow 1$ (as in GR). However we have to keep the following dimensionless quantity also fixed, $\rho:=\alpha/r_{h}^{2}$, whose necessity can be understood by observing the second term of the eqs. (172). This is somewhat peculiar, since it essentially requires scaling $G_{N}$ as we rescale the coordinates with the fluctuating boundary conformal factor. This problem originates from the fact that flat space lacks an additional dimensionful parameter, and $G_{N}$ is the only available quantity that comes in to play that role. This step can be justified by the fact that one recovers the same results from the flat limit ($\mathcal{P}\ell\rightarrow\infty$) of (A)dS, as shown in section 3. We remark, lastly, that our results remain unchanged whichever of the following three dimensionless parameters we keep fixed in the high-temperature limit: $\alpha/r_{h}^{2},\ \alpha/e^{\boldsymbol{\omega}}\mathfrak{r}r_{h}\ \text{or}\ \alpha/e^{2{\boldsymbol{\omega}}}\mathfrak{r}^{2}$. In the $y\rightarrow 1$ limit, eqs. (172) become:
@@
We now consider Einstein-Maxwell-Gauss-Bonnet gravity in flat space with line element ($\ell\rightarrow\infty,k=1$ in eq. (17)):
@@
Inverting these equations in this regime, we obtain:
@@ Appendix E Near the AdS Boundary in EMGB Gravity
In this appendix we discuss aspects of timelike boundaries in Einstein-Maxwell-Gauss-Bonnet gravity, placed near the conformal boundary of AdS. As uncovered by Fefferman and Graham (FG) FeffermanGraham1985 and generalized to arbitrary Lovelock metrics in ref. albin2020poincare , the metric of every asymptotically AdSd+1 spacetime admits an expansion near the conformal boundary of AdS as (notice that $\ell_{\textrm{\tiny AdS}}$ is the AdS curvature radius, which does not coincide with the $\ell$ appearing in the action for higher curvature gravity solutions):
@@
$\mathcal{P}\ell_{\textrm{\tiny AdS}}|_{\partial\textrm{\tiny AdS}}$ $=\mathcal{K}\ell_{\textrm{\tiny AdS}}|_{\partial\textrm{\tiny AdS}}+\frac{\alpha}{2(d-1)(d-2)}\mathcal{Q}\ell_{\textrm{\tiny AdS}}|_{\partial\textrm{\tiny AdS}}$ (184)
$=d\left(1-\frac{2}{3}\tilde{\alpha}g^{-2}(\tilde{\alpha})\right)+\frac{2\tilde{\alpha}\ell^{2}}{d-1}\tilde{R}(g^{(0)}_{ij})e^{-2\mathfrak{r}}+\mathcal{O}\left(e^{-2\mathfrak{r}}\right)\ .$
@@ Appendix F Entropy Universality in Lovelock Gravity, and Beyond
As emphasized in the main body of this work, entropy in Lanczos-Lovelock gravity is independent of the kind of boundary conditions one wishes to impose. Rather, it is an intrinsic feature of the horizon, a fact that can be linked to the Wald formula Wald:1993nt ; PhysRevLett.70.3684 ; Jacobson:1993vj ; Jacobson:1994qe ; Dong:2013qoa . However, within the framework of Euclidean thermodynamics there exists a simpler computation, see e.g., refs. Carlip:1993sa ; Banados:1993qp , demonstrating directly that the only contribution to the entropy comes exclusively from the horizon.
@@
where again $a_{d}$ is determined by the boundary conditions, so that the action (186) has a good variational problem. In fact, the precise form of the boundary action is not important for the following arguments. All that matters is that there exists a local boundary term that makes the action stationary under first order variations of the metric. We then want to compute the entropy though the usual formula:
@@
where $\beta$ is the proper boundary [^30] Or the conformal temperature $\tilde{\beta}$ used for (E)CBC or whichever temperature-like quantity defines the thermal ensemble. temperature. We are, therefore, interested in how the action changes to first order in $\beta$. To that extend, we can vary $\beta$ infinitesimally while keeping the bulk solution fixed. [^31] This is the opposite of what we usually do, where we compute the on-shell action for any $\beta$, appropriately adjusted so that we satisfy the equations of motion everywhere in the interior, and then differentiate to get the entropy. The two approaches coincide for first order variations of the metric, because the action (186) is stationary under those. Recall that, near the horizon $r_{h}$, metrics of the general form (15) become
@@
This derivation can also be readily generalized to arbitrary Lovelock gravity theories, for ensembles that are defined by boundary conditions that possess a stationary action under first order metric variations, possibly after the addition of a local boundary term. The only essential difference from before is the contribution of the bulk action at the position of the conical excess, which is just the sum of general Lovelock densities $\chi_{2n}(g)$ evaluated on the manifold $(\mathcal{M},g)$:
@@
with some dimensionful coefficients $\lambda_{n}$. For an opening angle of $\epsilon$ away from $2\pi$, the contribution of the density $\chi_{2n}(g)$ at the excess is given by (generalizing eq. (190)) Fursaev:1995ef :
@@
where the density $\chi_{2(n-1)}(\gamma)$ is evaluated for the codimension-$2$ horizon metric $\gamma$ with curvature quantities intrinsic to the horizon. The calculation proceeds as previously, yielding:
@@
Rather remarkably, this reproduces the Jacobson-Myers entropy Jacobson:1994qe , the generalization of the Wald entropy Wald:1993nt ; Iyer:1995kg . [^32] This is slightly ambiguous, since in the examples of Killing horizons considered here the additional extrinsic curvature terms vanish and the Jacobson-Myers entropy reduces to the Wald entropy. In the example of Einstein-Gauss-Bonnet gravity, we get:
@@
which is the familiar expression (45). Again, the discussion is unaltered by the addition of $U(1)$ gauge fields. A similar reasoning can also be followed for more general actions that are covariant functionals of the Riemann tensor, in the spirit of ref. Wald:1993nt .

### v2

@@ A.3 Variation of the Maxwell Action
In our conventions (9), the Lorentzian Maxwell action reads
$\mathcal{I}_{\text{M}}=-\frac{1}{64\pi G_{N}}\int_{\mathcal{M}}d^{d+1}x\sqrt{-g}F^{\mu\nu}F_{\mu\nu}\ .$ (146)
@@
$\mathcal{I}_{\text{M}}=-\frac{1}{32\pi G_{N}}\int_{\mathcal{M}}F\wedge\star F\ .$ (147)
@@
$\delta\mathcal{I}_{\text{M}}=-\frac{1}{16\pi G_{N}}\int_{\mathcal{M}}d\delta A\wedge\star F=-\frac{1}{16\pi G_{N}}\left(\epsilon\int_{\partial\mathcal{M}}\delta A\wedge\star F-\int_{\mathcal{M}}\delta A\wedge d\star F\right)\ ,$ (148)
@@
$\delta\mathcal{I}_{\text{M}}\hat{=-}\frac{\epsilon}{16\pi G_{N}}\int_{\partial\mathcal{M}}\delta A\wedge\star F\ ,$ (150)
@@
$\delta\mathcal{I}_{\text{M}}\hat{=}-\frac{\epsilon}{16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}A\wedge\star F\right)+\frac{\epsilon}{16\pi G_{N}}\int_{\partial\mathcal{M}}A\wedge\delta(\star F)\ .$ (151)
@@
$\mathcal{I}_{\text{M}}^{(\partial\mathcal{M})}=\frac{\epsilon}{16\pi G_{N}}\int_{\partial\mathcal{M}}A\wedge\star F\equiv\frac{\epsilon}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}F^{\mu\nu}n_{\mu}A_{\nu}\ ,$ (152)
then from the remaining term of eq. (151), the required boundary condition for a good variational problem is $\delta(\sqrt{|h|}F^{\mu\nu}n_{\mu})|_{\partial\mathcal{M}}=0$. This corresponds to fixing the electric charge, since
@@
From this derivation it is also evident that boundary conditions on the metric and the gauge field can be imposed independently.
@@ Appendix B Local and Covariant Boundary Terms for CBC
In this appendix we demonstrate that there does not exist a local and covariant boundary term which yields a well-defined variational principle for the bulk action (9) subject to the CBCs (3). For simplicity, we restrict to Euclidean signature.
To begin, note that after taking into account the Gauss-Codazzi equations, there exist nine independent local and covariant boundary terms one can write down, whose integrands take the form Bunch (1981) (again curvature tensors with overhead tilde are evaluated for the boundary metric)
@@
where $\mathcal{L}_{n}$ is the Lie derivative with respect to the normal $n$ to the boundary $\partial\mathcal{M}$. As noted e.g., in ref Bunch (1981), the variations of $D^{i}D_{i}\mathcal{K},\,D_{i}D_{j}\mathcal{K}^{ij}$ yield third order derivatives of the metric variation, and there is no linear combination of the two or some contribution from the variations of any of the other of the seven terms of eq. (154) that can cancel them. Hence, these cannot be included in the boundary term. A similar argument shows that the variations of the last two terms in eq. (154) cannot be canceled either.
@@
We would now like to examine whether there exists a combination of the coefficients $c_{i}$, such that the variation of eq. (155) cancels the boundary pieces of the on-shell variation of the bulk E(M)GB action after imposing CBC (3). Since, for CBC, we fix the conformal class of the induced metric, it will be convenient to decompose the various variations in terms of the variation of the conformal class and the variation of the Weyl factor. To accomplish that, we recall that for two $d$-dimensional metrics that are related via a Weyl rescaling $\tilde{h}_{ij}=\Omega^{2}(y^{m})h_{ij}$, the extrinsic curvature, Ricci tensor and Ricci scalar are related by (see e.g., appendix D in Wald (1984)),
$\begin{cases}\tilde{\mathcal{K}}_{ij}=\Omega\big(\mathcal{K}_{ij}+h_{ij}n^{k}D_{k}\log{\Omega}\big)\ ,\\ \tilde{R}_{ij}=R_{ij}-(d-2)D_{i}D_{j}\log{\Omega}-h_{ij}h^{pq}D_{p}D_{q}\log{\Omega}\\ \qquad\quad+(d-2)D_{i}\log{\Omega}D_{j}\log{\Omega}-(d-2)h_{ij}h^{pq}D_{p}\log{\Omega}D_{q}\log{\Omega}\ ,\\ \tilde{R}=\Omega^{-2}\Big(R-2(d-1)h^{ij}D_{i}D_{j}\log{\Omega}-(d-1)(d-2)h^{ij}D_{i}\log{\Omega}D_{j}\log{\Omega}\Big)\ ,\end{cases}$ (156)
@@
$+h^{-\frac{1}{d}}\mathcal{K}D^{i}\log{\sqrt{h}}D_{i}\log{\sqrt{h}}\big)\Bigg)\delta\sqrt{h}\Bigg]\,$,
@@ Appendix C On-Shell Actions for EMGB Solutions
In this appendix, we briefly describe how to compute the on-shell actions (9) for the solutions  (17). In computing the on-shell actions, the following curvature quantities for the general metric of the form (15) will be useful Deser and Ryzhov (2005); Haroon et al. (2020),
@@
where we defined $\lambda=d-1,\ a(r)=f^{\prime}(r)/2\sqrt{f(r)}\ \text{and}\ b(r)=\sqrt{f(r)}/r$. Moreover, the trace of the EGB field equations is (recall the definition of $\mathcal{E}_{\mu\nu}$ from eq. (116) and also that $\mathcal{L}_{\text{GB}}=R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}-4R^{\mu\nu}R_{\mu\nu}+R^{2}$),
@@
which has to be zero on-shell, and thus,
@@
Using this, we can re-express the bulk term of the action (9) as
@@
since the $r$ integral reduces to a boundary term by successive integrations by parts,
@@
A useful identity is given by (here $q=0$),
@@
Using eq. (164), the bulk on-shell action simplifies to,
@@
Therefore, the GB boundary term (118) can be collectively expressed as
@@
while the corresponding GR boundary term (96) reads,
@@
Putting eqs. (165), (167), (168) together we have the following full on-shell action (here we specialize to $k=a_{d}=b_{d}=1$ and $q=0$),
@@ Appendix D Flat EMGB black hole
We consider first the case of an asymptotically flat BD black hole with spherical horizon ($q=0,\Lambda=0,k=1$ in eq. (17)), with line element
@@
The boundary is at $r=e^{\boldsymbol{\omega}}\mathfrak{r}$ and we consider thermodynamics for the region between the black hole horizon at $r=r_{h}$ and the boundary $\partial\mathcal{M}$. As in (A)dS, inverting the boundary data eqs. (12) for the metric (170) to obtain $r_{h},e^{\boldsymbol{\omega}}\mathfrak{r}$ as a function of $\{\tilde{\beta},\mathcal{P}\}$ is not possible analytically, even for the simplest case of $d=4$. Nonetheless, perturbatively in the GB coupling, we get (recall that $y=r_{h}/e^{\boldsymbol{\omega}}\mathfrak{r}$),
@@
Due to the absence of additional scales in flat space, the high-temperature limit corresponds to sending $y\rightarrow 1$ (as in GR). However we have to keep the following dimensionless quantity also fixed, $\rho:=\alpha/r_{h}^{2}$, whose necessity can be understood by observing the second term of the eqs. (171). This is somewhat peculiar, since it essentially requires scaling $G_{N}$ as we rescale the coordinates with the fluctuating boundary conformal factor. This problem originates from the fact that flat space lacks an additional dimensionful parameter, and $G_{N}$ is the only available quantity that comes in to play that role. This step can be justified by the fact that one recovers the same results from the flat limit ($\mathcal{P}\ell\rightarrow\infty$) of (A)dS, as shown in section 3. We remark, lastly, that our results remain unchanged whichever of the following three dimensionless parameters we keep fixed in the high-temperature limit: $\alpha/r_{h}^{2},\ \alpha/e^{\boldsymbol{\omega}}\mathfrak{r}r_{h}\ \text{or}\ \alpha/e^{2{\boldsymbol{\omega}}}\mathfrak{r}^{2}$. In the $y\rightarrow 1$ limit, eqs. (171) become
@@
We now consider Einstein-Maxwell-Gauss-Bonnet gravity in flat space with line element ($\ell\rightarrow\infty,k=1$ in eq. (17)),
@@
Inverting these equations in this regime, we obtain,
@@ Appendix E Near the AdS Boundary in EMGB Gravity
In this appendix we discuss aspects of timelike boundaries in Einstein-Maxwell-Gauss-Bonnet gravity, placed near the conformal boundary of AdS. As uncovered by Fefferman and Graham (FG) Fefferman and Graham (1985) and generalized to arbitrary Lovelock metrics in ref. Albin (2020), the metric of every asymptotically AdSd+1 spacetime admits an expansion near the conformal boundary of AdS as (notice that $\ell_{\textrm{\tiny AdS}}$ is the AdS curvature radius, which does not coincide with the $\ell$ appearing in the action for higher curvature gravity solutions),
@@
$\mathcal{P}\ell_{\textrm{\tiny AdS}}|_{\partial\textrm{\tiny AdS}}$ $=\mathcal{K}\ell_{\textrm{\tiny AdS}}|_{\partial\textrm{\tiny AdS}}+\frac{\alpha}{2(d-1)(d-2)}\mathcal{Q}\ell_{\textrm{\tiny AdS}}|_{\partial\textrm{\tiny AdS}}\,$, (183)
$=d\left(1-\frac{2}{3}\tilde{\alpha}g^{-2}(\tilde{\alpha})\right)+\frac{2\tilde{\alpha}\ell^{2}}{d-1}\tilde{R}(g^{(0)}_{ij})e^{-2\mathfrak{r}}+\mathcal{O}\left(e^{-2\mathfrak{r}}\right)\ ,$
@@ Appendix F Entropy Universality in Lovelock Gravity, and Beyond
As emphasized in the main body of this work, entropy in Lanczos-Lovelock gravity is independent of the kind of boundary conditions one wishes to impose. Rather, it is an intrinsic feature of the horizon, a fact that can be linked to the Wald formula Wald (1993); Jacobson and Myers (1993); Jacobson et al. (1994b, a); Dong (2014). However, within the framework of Euclidean thermodynamics there exists a simpler computation, see e.g., refs. Carlip and Teitelboim (1995); Banados et al. (1994), demonstrating directly that the only contribution to the entropy comes exclusively from the horizon.
@@
where again $a_{d}$ is determined by the boundary conditions, so that the action (185) has a good variational problem. In fact, the precise form of the boundary action is not important for the following arguments. All that matters is that there exists a local boundary term that makes the action stationary under first order variations of the metric. We then want to compute the entropy though the usual formula,
@@
where $\beta$ is the proper boundary temperature. [^31] Or the conformal temperature $\tilde{\beta}$ used for (E)CBC or whichever temperature-like quantity defines the thermal ensemble. We are, therefore, interested in how the action changes to first order in $\beta$. To that extend, we can vary $\beta$ infinitesimally while keeping the bulk solution fixed. [^32] This is the opposite of what we usually do, where we compute the on-shell action for any $\beta$, appropriately adjusted so that we satisfy the equations of motion everywhere in the interior, and then differentiate to get the entropy. The two approaches coincide for first order variations of the metric, because the action (185) is stationary under those. Recall that, near the horizon $r_{h}$, metrics of the general form (15) become
@@
This derivation can also be readily generalized to arbitrary Lovelock gravity theories, for ensembles that are defined by boundary conditions that possess a stationary action under first order metric variations, possibly after the addition of a local boundary term. The only essential difference from before is the contribution of the bulk action at the position of the conical excess, which is just the sum of general Lovelock densities $\chi_{2n}(g)$ evaluated on the manifold $(\mathcal{M},g)$,
@@
with some dimensionful coefficients $\lambda_{n}$. For an opening angle of $\epsilon$ away from $2\pi$, the contribution of the density $\chi_{2n}(g)$ at the excess is given by (generalizing eq. (189)) Fursaev and Solodukhin (1995),
@@
where the density $\chi_{2(n-1)}(\gamma)$ is evaluated for the codimension-$2$ horizon metric $\gamma$ with curvature quantities intrinsic to the horizon. The calculation proceeds as previously, yielding
@@
Rather remarkably, this reproduces the Jacobson-Myers entropy Jacobson et al. (1994a), the generalization of the Wald entropy Wald (1993); Iyer and Wald (1995). [^33] This is slightly ambiguous, since in the examples of Killing horizons considered here the additional extrinsic curvature terms vanish and the Jacobson-Myers entropy reduces to the Wald entropy. For Einstein-Gauss-Bonnet gravity, we get
@@
which is the familiar expression (45). Again, the discussion is unaltered by the addition of $U(1)$ gauge fields. A similar reasoning can also be followed for more general actions that are covariant functionals of the Riemann tensor, in the spirit of ref. Wald (1993).